# Lessons from the high-resolution spectroscopy of AW UMa and $\epsilon\,$CrA: Is the Lucy model valid?

### v4

@@ 7.2 The Stȩpień model
The Stȩpień (2009) model of W UMa-type binaries focuses on a matter largery ignored in the Lucy model, namely, the solution for the energy flow between the components. This model predicts inter-binary velocities, thereby undermining the main assumption of the Roche equipotentials, which is central to the Lucy photometric model and the basis for photometric determinations of $q_{\rm ph}$. While Roche equipotentials lose their meaning, some regions of gravitational dominance will still exist for moderate internal velocities. We are using the term “Roche lobes” in this sense – as a convenient construct, similar to how it has already been employed in binary evolution models – but not as strict, well defined surfaces. The new model offers a wide range of solutions for W UMa-type binaries formed from differentially evolved components and observed as various subclasses of such systems (Stȩpień & Kiraga, 2013).
@@
The primary transfers its matter to the less-massive secondary, which is expected not to expand ecessively during the process. This suggests that the secondary is a low-mass star or the core of a mildly evolved MS star, possibly resulting from a previous binary mass exchange or a component exchange in a triple system. The stream from the primary is massive and forms an equatorial “bulge” on the secondary. Its equatorial extension is subject to the binary gravitational forces, allowing it to fill the entire critical lobe of the star. As observed more clearly in the case of $\epsilon\,$CrA (Figure 2), the flow appears to exhibit a range of rotational velocities similar to that of a Roche-model-filling secondary at phases around $\phi\simeq 0.25$, but it settles to a smaller velocity range at the $\phi\simeq 0.75$ half of the orbit. We note that the secondary component’s signature is very difficult to observe leading to a particularly large measurement scatter in Figures 2 and 7.
While the photospheric sound velocity for F-type stars, such as the AW UMa and $\epsilon\,$CrA primaries, is of the order 7 km s-1, the energy-carrying flow originates in sub-photospheric layers, where velocities may reach 30 – 50 km s-1. These velocities remain small relative to free-fall or orbital velocities, so gravitational forces still dominate in controlling the geometric properties of the flow. The flow continues around the secondary and returns to the primary. At all phases, the matter from primary component dominates the external view of the binary. The actual secondary star may be almost entirely hidden within the flow. The changing orbital aspect and eclipses of the flow are observed as the W UMa-type variability. The binary would be expected to look very differently at large inclinations, but in such cases, it would not be detected as a variable star.
@@ 8 Conclusions
Both binaries – particularly AW UMa – have been repeatedly observed photometrically, and their light curves perfectly agree with the common-equipotential light-curve-synthesis model of Lucy (1968a, b). However, the spectroscopic view presents a more complex picture: the binaries are not rotating as solid bodies, as required by the strict Lucy model, but instead exhibit several phenomena related to circumbinary flow with velocities not exceeding 30 – 50 km s-1. Velocities of that order, which are moderate compared to the free-fall or orbital velocities of about 300 – 350 km s-1, are predicted in the Stȩpień (2009) model of W UMa binaries. This model envisages the energy transport between the binary components as an optically thick stream or belt of matter originating from the Roche-lobe overfilling primary component (Section 7.2). The observed W UMa-type variations are due to the changing visibility and eclipses of this energy-carrying belt.
The velocity field around the secondary component carries important information relevant to the Stȩpień (2009) model. We note (Figure 2) that the observed rotational velocities of the $\epsilon\,$CrA secondary are very similar to those expected for a rigidly rotating Roche model in the first half of the orbit ($\phi=0.25$). In contrast, the second half of the orbit ($\phi=0.75$) shows possible settling of the velocities to lower rotational velocities. This behavior is exactly as predicted for a flow that initailly fills the critical lobe and then loses its energy, drifting closer to the secondary’s otherwise invisible surface. The same effect is not as clearly visible in the much more complicated case of AW UMa (Figure 7), which was also a less-well observed binary.
@@
The tendency for several ESPs in AW UMa does not have an explanation. Being less stable than the single, persistent ESP visible in $\epsilon\,$CrA, the ESPs showed a tendency to evolve within hours. Their variability may indicate that the flow around the AW UMa secondary, as envisaged in the Stȩpień model, is more erratic than that in $\epsilon\,$CrA or is subject to a localized pileup. We note the opposite tendencies in the systematic orbital-period changes of the two binaries (Section 2): As indicated by the sign of the period changes, the net mass exchange between the components of AW UMa appears to be directed from the more massive primary to the less massive secondary, while it is in the opposite direction in $\epsilon\,$CrA. We may speculate that the multitude of ESPs on the AW UMa secondary signals some sort of “overload” of the flow due to the abundance of the incoming matter at this particular stage of binary evolution.
@@
By explaining the energy-transfer problem in W UMa-type binaries and providing reasons for the $q_{\rm ph}<q_{\rm sp}$ tendency, the Stȩpień model casts doubt on the applicability of the common-equipotential Lucy model for deriving physical data for these binaries. Yet, the new model does not provide any predictions about the size of the deviations from the Lucy model, nor does it suggest ways how to correct $q_{\rm ph}$ results. Are the results already obtained with the photometric Lucy model still useful? This question applies to large-scale projects utilizing state-of-the-art light-curve synthesis codes (Prša et al., 2016; Conroy et al., 2020; Wilson et al., 2020) for the currently ongoing large photometric surveys, following the example of earlier explorations based on ASAS (Pilecki, 2009), OGLE (Soszyński et al., 2015) surveys, or to the large compilation of almost 700 previous results by Latković et al. (2021). A convenient tool for massive handling of statistical $q_{\rm ph}$ data – implicitly assuming a strict acceptance of the Roche model – has already been developed by Pešta & Pejcha (2023). Another question is: Are summaries utilizing mixed $q_{\rm ph}$ and $q_{\rm sp}$ data, such as those by Gazeas & Stȩpień (2008), Latković et al. (2021) or Gazeas (2024), giving us an unbiased picture?
While the Roche-model “short cut” to massive determination of $q$ appears to posses a serious difficulty in providing biased results, we currently have no reasons to expect any similar bias in spectroscopic results. The case of AW UMa suggests that medium-resolution ($R\simeq 15,\!000-20,\!000$) data may be adequate for an acceptable determination of $q_{\rm sp}$, but that analysis of detailed profiles and detection of ESPs require a higher resolution ($R>30,\!000$). The binary FG Hya ($V\simeq 10$, $q_{\rm sp}=0.14$) is the only case of the inverted mass-ratio discrepancy (Section 5, Figure 8), and it urgently requires a new spectroscopic determination. Binaries with less extreme mass ratios particularly need spectroscopic work: we already have tantalizing indications (Pribulla et al., 2006; Pribulla & Rucinski, 2008) that medium-resolution RV profiles for V566 Oph ($V\simeq 7.5$, $q_{\rm sp}=0.263\pm 0.012$) look – unlike those of our two binaries – as predicted by the Lucy model. Interpreted within the Stȩpień model, this may indicate that the circumbinary flow fully dominates the spectroscopic profile shape for V566 Oph, in contrast to our two, very low mass-ratio systems where the primaries are dominant. It is possible that the flow dominance is a typical situation for W UMa binaries with mass ratios larger than some specific value. We note that the two binaries considered here are not cases of extremely small $q_{\rm ph}$, as smaller values have already been reported in the literature (Wadhwa et al., 2021; Li et al., 2022; Guo et al., 2023).
@@
For a final summary: The simplistic Roche-model light-curve calculations have not significantly advanced our understanding of W UMa binaries. While spectroscopy is essential, it remains difficult and costly. We have some insight into how these binaries form in small stellar groups and that they likely consist of differentially evolved components. However, we are still in the early stages of understanding the structure and surprising stability of these binaries. The roles of tidal energy and angular momentum dissipation remains to be studied and incorporated into models similar to those of Oka et al. (2002). Such models could help in interpreting spectroscopic results, partly regarding the complex velocity fields observed within these binaries.
I would like to thank Staszek Zoła and Michał Siwak for useful suggestions and comments. Special thanks go to Kazik Stȩpień, the originator of the new model for W UMa binaries for his insightful ideas, interesting discussions, and suggestions on the presentation of arguments.

### v5

@@ 7.2 The Stȩpień model
The Stȩpień (2009) model of W UMa-type binaries focuses on an aspect largely ignored in the Lucy model: the solution for the energy flow between the components with a particular attention to the Coriolis force, which organizes the flow and makes it concentrated to the equatorial regions. The model predicts inter-binary velocities so that the Roche equipotentials cannot be defined. However, for sufficiently small velocities, gravity and centrifugal forces dominate, restoring the geometrical properties often described by the term of the “Roche lobes”. The Stȩpień model offers a wide range of solutions for W UMa-type binaries formed from differentially evolved components and observed as various subclasses of such systems (Stȩpień & Kiraga, 2013).
@@
The primary transfers its matter to the less-massive secondary, which is expected not to expand excessively during the process. This suggests that the secondary is a low-mass star or the core of a mildly evolved MS star, possibly resulting from a previous binary mass exchange or a component exchange in a triple system. The stream from the primary is massive and forms an equatorial “bulge” on the secondary. Its equatorial extension is subject to the binary gravitational forces, allowing it to fill the entire critical lobe of the star. As observed more clearly in the case of $\epsilon\,$CrA (Figure 2), the flow appears to exhibit a range of rotational velocities similar to that of a Roche-model-filling secondary at phases around $\phi\simeq 0.25$, but it settles to a smaller velocity range at the $\phi\simeq 0.75$ half of the orbit. We note that the secondary component’s signature is very difficult to observe leading to a particularly large measurement scatter in Figures 2 and 7.
While the photospheric sound velocity for F-type stars, such as the AW UMa and $\epsilon\,$CrA primaries, is of the order 7 km s-1, the energy-carrying flow originates in sub-photospheric layers, where velocities may reach 30 – 50 km s-1. These velocities remain small relative to free-fall or orbital velocities, so gravitational forces still dominate in controlling the geometric properties of the flow and the overall geometry does not deviate appreciably from the Roche model. The flow is confined to the equatorial regions. It continues around the secondary and returns to the primary. At all phases, the matter from primary component dominates the external view of the binary. The actual secondary star may be almost entirely hidden within the flow. The changing orbital aspect and eclipses of the flow are observed as the W UMa-type variability. The binary would be expected to look very differently at large inclinations, but in such cases, it would not be detected as a variable star.
@@ 8 Conclusions
Both binaries – particularly AW UMa – have been repeatedly observed photometrically, and their light curves perfectly agree with the common-equipotential light-curve-synthesis model of Lucy (1968a, b). However, the spectroscopic perspective presents a more complex picture of a circumbinary flow with velocities not exceeding 30 – 50 km s-1. Such velocities, moderate compared to the free-fall or orbital velocities of about 300 – 350 km s-1, are predicted in the Stȩpień (2009) energy transport model of W UMa binaries. This model envisages an optically thick stream or belt of matter originating from the Roche-lobe-overfilling primary component, which retains its heat during the transport (Section 7.2). The observed W UMa-type variations result from the changing visibility and eclipses of this energy-carrying belt, whose eclipse and visibility-aspect variations closely resemble those of the Roche model geometry.
We note (Figure 2) that the observed rotational velocities of the $\epsilon\,$CrA secondary are very similar to those expected for a rigidly rotating Roche model in the first half of the orbit ($\phi=0.25$). In contrast, the second half of the orbit ($\phi=0.75$) shows possible settling of the velocities to lower rotational velocities. This behavior is exactly as predicted for a flow that initailly fills the critical lobe and then loses its energy, drifting closer to the secondary’s otherwise invisible surface. The same effect is not as clearly visible in the much more complicated case of AW UMa (Figure 7), which was also a less-well observed binary.
@@
The difference between $\epsilon\,$CrA and AW UMa in the number of ESPs currently lacks an explanation. While the ESPs are most likely driven by local thermodynamical instabilities and are chaotic in nature, it is possible that the two binaries differ in the type or extent of the flow. The stronger ESP variability in AW UMa may indicate that the flow around its secondary, as envisaged in the Stȩpień model, is more erratic than that in $\epsilon\,$CrA or is subject to a localized pileup. We also note opposite tendencies in the systematic orbital-period changes of the two binaries (Section 2):
Based on the sign of the period changes, the net mass exchange between the components of AW UMa appears to be directed from the more massive primary to the less massive secondary, whereas in $\epsilon\,$CrA, it flows in the opposite direction. One possible explanation is that the multitude of ESPs on the AW UMa secondary signals an ”overload” of the flow, caused by the abundance of incoming matter at this particular stage of binary evolution
@@
While the Stȩpień model may suggest directions for exploring the $q_{\rm ph}<q_{\rm sp}$ tendency, it does not provide predictions regarding the magnitude of deviations from the Lucy model or suggest methods to reconcile $q_{\rm ph}$ or $q_{\rm sp}$ results. Currently, the photometric approach dominates, utilizing complex, state-of-the-art light-curve synthesis codes (Prša et al., 2016; Conroy et al., 2020; Wilson et al., 2020) which have largely superseded earlier explorations based on based on ASAS (Pilecki, 2009), OGLE (Soszyński et al., 2015) surveys, and the extensive compilation of nearly 700 results by Latković et al. (2021). A convenient tool for a large-scale handling of statistical $q_{\rm ph}$ data – implicitly assuming a strict acceptance of the Roche model – has already been developed by Pešta & Pejcha (2023). But do summaries combining $q_{\rm ph}$ and $q_{\rm sp}$ data, such as those by Gazeas & Stȩpień (2008), Latković et al. (2021) or Gazeas (2024), provide an unbiased picture?
New spectroscopic investigations of W UMa-type binaries are urgently needed. Observations of AW UMa at two spectral resolutions suggests that medium-resolution ($R\simeq 15,\!000-20,\!000$) data may suffice for an acceptable determination of $q_{\rm sp}$. However, detailed analysis of spectral profiles and detection of ESPs require a higher resolution ($R>30,\!000$). The binary FG Hya ($V\simeq 10$, $q_{\rm sp}=0.14$) is the only case of the inverted mass-ratio discrepancy (Section 5, Figure 8), and it urgently requires a new spectroscopic determination. Binaries with less extreme mass ratios particularly need spectroscopic work: we already have tantalizing indications (Pribulla et al., 2006; Pribulla & Rucinski, 2008) that medium-resolution RV profiles for V566 Oph ($V\simeq 7.5$, $q_{\rm sp}=0.263\pm 0.012$) look – unlike those of our two binaries – as predicted by the Lucy model. Interpreted within the Stȩpień model, this may indicate that the circumbinary flow fully dominates the spectroscopic profile shape for V566 Oph, in contrast to our two, very low mass-ratio systems where the primaries are dominant. It is possible that the flow dominance is a typical situation for W UMa binaries with mass ratios larger than some specific value. We note that the two binaries considered here are not cases of extremely small $q_{\rm ph}$, as smaller values have already been reported in the literature (Wadhwa et al., 2021; Li et al., 2022; Guo et al., 2023).
@@
For a final summary: While spectroscopy is essential for our understanding of W UMa binaries, it remains difficult and costly. We have some insight into how these binaries form in small stellar groups and that they likely consist of differentially evolved components. However, we are still in the early stages of understanding the structure and surprising stability of these binaries. The roles of tidal energy and angular momentum dissipation remains to be studied and incorporated into models similar to those of Oka et al. (2002). Such models could help in interpreting spectroscopic results, partly regarding the complex velocity fields observed within these binaries.
I would like to thank Dr. Andrej Prša who, acting as a reviewer of the paper, suggested importnat improvements and corrections to the text. I would like to thank Staszek Zoła and Michał Siwak for useful suggestions and comments. Special thanks go to Kazik Stȩpień, the originator of the new model for W UMa binaries for his insightful ideas, interesting discussions, and suggestions on the presentation of arguments.